# On the Use of Large Language Models in Mutation Testing

### v2

@@ 5.1.2. Impact of Different Few-Shot Examples
Real Bug Det. 0.917 0.933 0.917 0.917 0.900
Coupling Rate 0.136 0.128 0.132 0.121 0.139
Ochiai Coefficient 0.467 0.450 0.434 0.434 0.367
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples.
@@
We further compute their similarity, shown as the lower part of Table 10. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude the performance of different prompts is very similar.
@@ 5.1.3. Performance of Using the Same Number of Mutations
In Section 4, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 16,785. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 12. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
@@
Metric GPT-3.5 GPT-4o GPT-4o-M DC-236b SC-16b CL-13b LEAM $\mu$Bert Major PIT
Comp. Rate 62.9% 77.4% 70.1% 73.0% 39.3% 74.2% 38.9% 18.5% 96.5% ——
Useless Mut. Rate 11.0% 7.3% 7.0% 8.9% 27.1% 39.0% 1.4% 1.4% 0.0% ——
Real Bug Det. 89.2% 90.8% 91.3% 89.7% 58.0% 73.5% 67.5% 55.0% 83.3% 43.7%
Coupling Rate 13.3% 14.6% 12.6% 13.4% 11.9% 10.5% 13.2% 13.2% 12.8% 2.4%
Ochiai Coefficient 35.8% 45.0% 40.8% 35.3% 15.3% 17.8% 17.0% 15.0% 34.1% 19.2%
5.2. Economic Costs
Closed-source LLMs (e.g., GPT LLMs) are typically billed based on the number of tokens used. As these models continue to evolve, their pricing rates also change, and we here discuss the costs of the proprietary models used in our experiments. For the three models involved, GPT-3.5 incurred a total cost of \$108, GPT-4o cost \$1531, and GPT-4o-Mini cost \$47. Therefore, for every 1,000 mutations, GPT-3.5, GPT-4o, and GPT-4o-Mini cost about \$0.16, \$2.84, and \$0.08, respectively. GPT-4o is significantly more expansive than the others. Given the effectiveness of mutations generated by GPTs, the cost is feasible for large-scale applications.
@@ 5.3. Implications
Based on our findings, we discuss the pros and cons of using LLMs for generating mutations and suggest potential improvements.
First, our study shows that existing LLMs have the potential to generate high-quality mutations for mutation testing with acceptable costs. The greatest advantage of LLMs is their ability to generate mutations that more closely mimic real bugs in behavior. For example, Listing 1 shows two compilable mutations generated by GPT-3.5 in Chart-1, which are theoretically beyond the capabilities of existing approaches. In Mutation-1, the LLM inferences there must exist leftBlock because of the occurrence of rightBlock, and replaces the variable correspondingly. This mutation is not killed by the bug-triggering test. Without natural language understanding capabilities (i.e., understanding the correspondence between “right” and “left”), it is difficult to create such mutations. In Mutation-2, the LLM modifies the else branch to else if, employing an integer variable seriesCount from context to synthesize the condition, showing the capabilities of code structural changing that existing methods cannot achieve. Also, this mutation is alive and can guide test enhancement.
@@
Second, our study explores how different prompts, models, contexts, and few-shot examples impact mutation generation performance. Our study highlights the importance of using proper prompts and models with stronger coding capabilities. In particular, designing prompts requires providing essential context information (such as the whole method), but including too much information (such as the target unit test) may reduce effectiveness.
Third, although LLMs achieve promising results in mutation testing, our study indicates that there is still significant room for improvement. Specifically, current LLMs generate a substantial number of mutations that are not compilable. Therefore, we analyze the types of mutations that lead to non-compilable code and identify some root causes of the LLMs. We advocate for more research efforts in program analysis and repair approaches to address error-prone code elements for mutation such as method invocations.
5.4. Threats to Validity
@@
The selected LLMs, programming language, datasets, and baseline approaches could be a threat to the validity of our results. To mitigate this threat, we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek), the most popular language (i.e., Java), and the most popular dataset Defects4J. We also employ state-of-the-art mutation testing approaches as baselines, including learning-based and rule-based.
@@
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (469), GPT-4o (290), GPT-4o-Mini (528), DeepSeek (318), StarChat (122), CodeLlama (236), LEAM (590), $\mu$Bert (66), and Major (217), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), $\mu$Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of the exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar.
@@ 6. Conclusion
In this paper, we systematically investigate the performance of LLMs in generating mutations. We evaluate their utility from several aspects and find LLMs have the advantage of generating diverse mutations that mimic the behaviour of real bugs. We also analyze and discuss some limitations of LLMs and point out further directions. We advocate research efforts to promote LLM for mutation testing in the future.
7. Data Availability
All the experimental data, code, and full results (including generated and manually labeled mutations) are publicly available (our, 2024).

### v3

@@ 4.1.3. Impact of Different Few-Shot Examples
QB-N: Select N Examples from QuixBug. QB-6a is the set of default examples, while QB-6b involves 6 other examples. D4J-6 involves random 6 examples from Defects4J.
@@
We further compute their similarity, shown as the bottom part of Table 14. We can find that their similarity exceeds 0.95, significantly exceeding the typical threshold of 0.85, and the $p$-values are all below 0.05. Therefore, we can conclude that although different few-shot examples have perturbation, their performance is very similar.
@@ 4.1.4. Performance of Using the Same Number of Mutations
In Section 3, we did not restrict the number of mutations generated by each approach. To study the performance of each approach under fixed quantity conditions, we followed the setting of existing study (Tian et al., 2022; Hariri et al., 2019) and limited the number of mutations generated by all methods to the minimum produced by $\mu$Bert (Degiovanni and Papadakis, 2022), which is 11397. For counts exceeding this number, we randomly selected the specified number of mutations for analysis. We conducted 10 rounds of random sampling comparisons, and the results are presented in Table 16. Therefore, under fixed quantity conditions, the mutations generated by two GPT-4 LLMs are still the closest to real bugs in behavior.
@@
Metric PIT Major LEAM $\mu$BERT GPT-3.5 GPT-4o GPT-4o-M DS-236b SC-16b CL-13b
Real Bug Det. 31.6% 68.8% 64.6% 66.0% 82.2% 87.4% 88.6% 87.0% 33.9% 68.5%
Coupling Rate 17.4% 45.4% 37.2% 43.6% 45.6% 48.8% 47.9% 47.9% 27.1% 43.6%
Ochiai Coefficient 12.6% 22.2% 14.1% 19.0% 26.5% 36.0% 31.8% 34.1% 15.4% 16.9%
Comp. Rate 96.9% 54.1% 26.1% 61.2% 75.5% 74.3% 74.6% 12.4% 69.5%
Dup Mut. Rate 0.0% 2.6% 1.9% 9.9% 7.7% 7.1% 7.6% 9.6% 38.3%
@@ 4.2. Implications
Based on our findings, we discuss the pros and cons of traditional and LLM-based mutation generation approaches, and suggest potential improvements.
4.2.1. Advantages of LLM-based Approaches
Our study demonstrates that LLMs can generate high-quality mutations for mutation testing at an acceptable cost. The key advantage of LLMs lies in their ability to produce mutations with stronger coupling relationships with real-world bugs, which is crucial for improving test effectiveness.
Additionally, LLMs could generate syntactically diverse mutations compared with traditional approaches, which is potentially the reason for generating high-coupling mutations. For example, Listing 1 presents two compilable mutations generated by GPT-3.5 for Chart-1, which are theoretically beyond the capabilities of traditional approaches. In Mutation-1, the LLM infers a relationship between variables (rightBlock implies the existence of leftBlock) and modifies the variable accordingly. This mutation is not killed by the original bug-triggering test, highlighting its potential for guiding test enhancement. Without natural language understanding (e.g., recognizing the correspondence between “right” and “left”), traditional approaches, both rule-based approaches and the approaches based on small-scale models, struggle to generate such context-aware mutations. In Mutation-2, the LLM transforms an else branch into an else if condition by synthesizing a logical condition using the integer variable seriesCount from the context, demonstrating an ability to introduce structurally meaningful changes that traditional approaches cannot achieve.
@@ 4.2.1. Advantages of LLM-based Approaches
Additionally, our study explores how different prompt designs, model choices, and contextual information influence mutation generation performance. We find that selecting models with stronger code generation capabilities (e.g., GPT-4o) and designing prompts carefully are critical to achieving optimal results.
4.2.2. Advantages of Traditional Approaches
Despite the advantages of LLM-generated mutations, traditional approaches still exhibit strengths in reliability and efficiency. They produce fewer non-compilable and duplicate mutations, with PIT achieving a 100% compilability rate and a 0% duplicate mutation rate, while Major follows closely with a 97.6% compilability rate and 0% duplicate mutation rate. In contrast, LLM-based approaches exhibit significantly lower compilability and higher duplicate mutation rates. Additionally, overall, traditional rule-based approaches generate a slightly lower equivalent mutation rate, further avoiding redundant costs. This suggests that traditional approaches remain highly practical when strict compilability and low redundancy are prioritized.
4.2.3. Challenges and Future Directions
Although LLMs demonstrate superior behavioral similarity to real bugs, our findings indicate substantial room for improvement. One major limitation is the high rate of non-compilable mutations, which limits their practical application in mutation testing. Our analysis reveals that common causes include syntactic errors in generated expressions, incorrect method invocations, and type mismatches. To address these challenges, future research should explore integrating LLM-based approaches with program analysis and repair techniques to refine mutation generation. Specifically, leveraging static analysis to filter or adjust syntactically incorrect mutations before execution can enhance their reliability.
5. Threats to Validity
In this section, we illustrate the threats to the validity of our study.
@@ 5. Threats to Validity
The selected programming language, mutation generation approaches, datasets, and LLMs could be a threat to the validity of our results. To mitigate this threat, employ the most widely studied mutation testing approaches as baselines, including learning-based and rule-based, the most widely used language (i.e., Java), and the most widely used dataset Defects4J, and we adopt the widely studied models covering a range of base models (i.e., GPT, StarCoder, CodeLlama, and DeepSeek).
@@
Another validity threat may be due to data leakage, i.e., the fact that the data in Defects4J (Just et al., 2014a) may be covered in the training set of the studied LLMs. To mitigate this threat, we employed another dataset ConDefects (Wu et al., 2024) which includes programs and faults that were made after the release time of the LLMs we use and thus have limited data leakage risk. Additionally, to increase confidence in our results, we also checked whether the tools can introduce exact matches (syntactically) with the studied faults. We hypothesize that in case the tools have been tuned based on specific fault instances, the tools would introduce at least one mutation that is an exact match with the faults we investigate. On Defects4J, the results are GPT-3.5 (133), GPT-4o (63), GPT-4o-Mini (138), DeepSeek (66), StarChat (13), CodeLlama (28), LEAM (287), $\mu$Bert (12), and Major (22), respectively. On ConDefects, the results are GPT-3.5 (22), GPT-4o (25), GPT-4o-Mini (31), DeepSeek (32), StarChat (35), CodeLlama (56), LEAM (56), $\mu$Bert (5), and Major (92), respectively. The comparison indicates that on Defects4J, GPTs and LEAM tend to produce significantly more exact matches than the other approaches. Interestingly, Major produces a similar number of exact matches with Codellama. $\mu$Bert yields significantly the fewest number of exact matches, indicating a minimal or no advantage for all these approaches (except from GPTs and LEAM) due to exact matches (in the case of Defects4J). Perhaps more interestingly, on the ConDefects dataset, which has not been seen by any of the tools, Major has the highest number of exact matches, indicating a minor influence of any data leakage on the reported results. Nevertheless, the LLMs we studied exhibit the same trend on the two datasets, achieving the Spearman coefficient of 0.943 and the Pearson correlation of 0.944, both with $p$-value less than 0.05, indicating their performance is similar on the two datasets.
The different experimental settings may also threaten the validity of our results. To address the threat, we explore the impacts of prompts, context length, few-shot examples, and mutation numbers on the performance of LLMs. The results show that different settings are highly similar. One potential threat to validity is that the metrics used in this study may not fully reflect the actual bug detection in practice. Our evaluation is conducted on the fixed versions of programs, rather than the faulty versions where real-world testing would typically be applied, which may introduce potential differences. In other words, our study operates under the clean program assumption, which may not fully capture the dynamics of bug detection in practical scenarios. However, to mitigate this threat, we use the widely adopted Defects4J 2.0 dataset, which consists of real bugs from well-maintained open-source projects, ensuring the relevance of our findings to real bugs,
@@
6. Related Work
In this section, we introduce the related work of our study.
6.1. Mutation Testing and Mutation Generation
Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015; Wang et al., 2025), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
Traditional mutation testing approaches generate mutations by pre-defined program transformation rules (i.e., mutation operators), which are referred to rule-based. For example, one widely used mutation operator is to change one arithmetic binary operator with another one (e.g., $a+b\mapsto a-b$). During the last two decades, the majority of existing mutation testing approaches have been rule-based (Jia and Harman, 2010; Amalfitano et al., 2022). Researchers primarily focus on approaches in Java (e.g., PIT (Coles et al., 2016), Major (Just et al., 2011), JavaLanche (Schuler and Zeller, 2009), MuJava (Ma et al., 2005), IBIR (Khanfir et al., 2023), etc.) and C/C++ (e.g., Proteum (Delamaro et al., 1996), Milu (Jia and Harman, 2008), WinMut (Wang et al., 2021), etc.), leveraging the mature ecosystems and tooling chains of these languages for mutation testing.
Recently, with the emergence of machine learning techniques, researchers proposed using deep learning models to generate mutations. For instance, DeepMutation (Tufano et al., 2019) utilizes machine sequence-to-sequence neural translation methods, while LEAM (Tian et al., 2022) employs learning-guided rule-based expansion. However, these approaches suffer from the challenges in generating syntactic correct mutations and mutations with diverse forms and often fail to generate syntactically valid mutations. Perhaps the closest to our work is that of Degiovanni et al. who proposed $\mu$Bert (Degiovanni and Papadakis, 2022). $\mu$Bert generates mutation by masking the code elements to be mutated and then predicting the mutation via Bert (Devlin et al., 2018).
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches. Dakhel et al.(Dakhel et al., 2024) and Ibrahimzada et al.(Ibrahimzada et al., 2023) also proposed generating artificial bugs using LLMs. These approaches mainly evaluate from a certain perspective, such as mutation score or program repair. Compared with these approaches, our study is more comprehensive and evaluates existing mutation generation approaches from many aspects. Specifically, we measure the behavioral similarity between real bugs and mutations using established metrics, evaluate the syntactic diversity of mutations, and analyze the non-compilable mutations produced by LLMs. Our findings illustrate both the strengths and limitations of LLMs for mutation generation, and highlight directions for future improvement.
6.2. LLMs for Code
Recently, Large Language Models (LLMs) have shown remarkable potential in a wide range of natural language processing (NLP) (Brown et al., 2020) and coding tasks (Wang et al., 2024; Fan et al., 2023). Existing LLMs are mostly based on Transformer (Vaswani et al., 2017) and are trained by a large corpus of data, including texts and source codes. LLMs are driven by instructions in natural languages, i.e., prompts, which guide the models to perform target tasks and are crucial to performance. The design of prompts plays a crucial role in obtaining accurate and useful responses from these models.
@@ 7. Conclusion
In this paper, we systematically investigate the performance of existing mutation generation appraoches across multiple dimensions, including their effectiveness, validity, and efficiency. Our study conducts extensive and detailed comparative experiments, covering both widely studied rule-based mutation generation approaches and LLM-based methods. Our evaluation results highlight that LLMs significantly outperform traditional approaches in generating diverse mutations that closely mimic real bugs. However, the results also expose key limitations of LLMs, particularly their tendency to generate a high proportion of non-compilable mutations. Therefore, we analyze the error types of the LLM-generated non-compilable mutations and identify member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations. Based on our findings, we emphasize the need for further research to enhance LLM-based mutation testing, addressing its current shortcomings while leveraging its strengths.